what is teh slope of 2x+y=3
Answer by jim_thompson5910(35256)

Solve for y:


2x+y=3


y=3-2x


y = -2x+3


The equation is now in y = mx+b form where m is the slope. Since m = -2, the slope is -2.


If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim